Derive van der waals eqiation from the ideal gas equation.
1
Expert's answer
2019-09-12T03:31:57-0400

Ideal gas equation PV=RT (n=1mol)

For real gass

1) We put the volume available to the molecules (Vm-b), where Vm is the molar volume , b is the volume that the molecules occupy."P(V_m-b)=RT"

2) We introduce a correction that takes into account the interaction between the molecules."P=p+\frac{a}{V_m^2}"

"(P+\frac{a}{V_m^2})(V_m-b)=RT"

https://en.m.wikipedia.org/wiki/Van_der_Waals_equation